OpenStudy (anonymous):

Find the integral of 1/(sqrt(t^3+2)) dt from 0 to x.

OpenStudy (anonymous):

\[\int_0^x\frac{1}{\sqrt{t^3+2}}dt\]?

OpenStudy (anonymous):

you are not going to find a nice closed form for this what is the actual question?

OpenStudy (anonymous):

Yes. Please continue.

OpenStudy (anonymous):

what does the question say? my guess: "find the derivative" but i could be wrong

OpenStudy (anonymous):

it cannot really say "find the integral" because it is an integral

OpenStudy (anonymous):

It's a multiple choice question. If f(x)=integral of 1/(sqrt(t^3+2)) dt from 0 to x, which of the following is FALSE? a) f(0)=0 b) f is continuous at x for all x greater than or equal to 0. c) f(1)>0 d) f'(1)=1/sqrt(3) e) f(-1)>0

OpenStudy (anonymous):

ok now we have a question

OpenStudy (anonymous):

\[f(x)=\int_0^x\frac{1}{\sqrt{t^3+2}}dt\] \[f(0)=\int_0^0\frac{1}{\sqrt{t^3+2}}dt=0\]

OpenStudy (anonymous):

Why does f(0)=0?

OpenStudy (anonymous):

so first one is right also \(f(x)=\int_0^x\frac{1}{\sqrt{t^3+2}}dt\) is a continuous function, so second one is right

OpenStudy (anonymous):

\[f(0)=\int_0^0\frac{1}{\sqrt{t^3+2}}dt\] yes?

OpenStudy (anonymous):

you are integrating over no path, or a path with length zero, so the integral is 0

OpenStudy (anonymous):

Okay, continue.

OpenStudy (anonymous):

\[f(x)=\int_0^x\frac{1}{\sqrt{t^3+2}}dt\] \[f(1)=\int_0^1\frac{1}{\sqrt{t^3+2}}dt\] and this is positive because the integrand is positive on this interval

OpenStudy (anonymous):

And how is f(1) positive?

OpenStudy (anonymous):

because the integrand is positive on the interval \((0,1)\)

OpenStudy (anonymous):

You mean 1/(sqrt(t^3+2)) is positive on (0, 1)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, continue.

OpenStudy (anonymous):

\[f(x)=\int_0^x\frac{1}{\sqrt{t^3+2}}dt\] \[f'(x)=\frac{1}{\sqrt{x^3+2}}\] and so \[f'(1)=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

Okay, continue?

OpenStudy (anonymous):

i guess the last one has to be wrong

OpenStudy (anonymous):

since the intgrand is positive on \((-1,0)\) then \[f(-1)=\int_0^{-1}\frac{1}{\sqrt{t^3+2}}dt\] must be negative

OpenStudy (anonymous):

because you are going backwards, from \(0\) to \(-1\) this \[\int_{-1}^0\frac{1}{\sqrt{t^3+2}}dt\] would be positive

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw